NAOJ GW Elog Logbook 3.2

Michael, Yuhang
After installing RF amplifiers and USB switches, we worked on the recovery today.
1. We reboot standalone to solve the problem of timing
2. We delete some second trend and released 10% space (this corresponds to about 1 month data)
3. We checked PLL p-pol frequency changed to 265MHz (53) with OPO 7.162 temperature and no green. It was 240MHz (48). In this case, BAB power before PBS is 0.46mW.
4. The mode matching from BAB into filter cavity was checked to be better than 93% (520 counts was observed for FC BAB tra)
5. Then we found main laser turned off by itself. We didn't notice how it happened. Then we checked dataviewer. The PR/BS/INPUT/END oplevs, FC GR/IR tra, FC GR correction/error were checked. Figure 1 shows this check. From this check, we could find
- The first change happend for FC GR/IR tra, FC GR error suddenly.
- PR oplev didn't have any change. BS oplev changed gradually due to pointing loop. INPUT oplev didn't change because of its wrong set-up. END oplev didn't have sudden change, either. So this main laser sudden off should not come from suspension.
- FC GR correction looks not that sudden compared with GR tra. So this problem seems also not come from a sudden large laser frequency or cavity length change.
So it is still not for sure why we see this change.